You are on a merry go round with radius 6.1 m. What is the centripetal force on you if your velocity is 2.7 m/s and you weigh 7.5 kilograms?

1
Expert's answer
2022-04-04T17:09:36-0400
"F_c=\frac{mv^2}{r}"

"F_c=\frac{7.5*2.7^2}{6.1}=9.0\:\rm N"